# Parameters vs FLOPs: Scaling Laws for Optimal Sparsity for Mixture-of-Experts Language Models

### v1

@@ Abstract
Scaling the capacity of language models has consistently proven to be a reliable approach for improving performance and unlocking new capabilities. Capacity can be primarily defined by two dimensions: the number of model parameters and the compute per example. While scaling typically involves increasing both, the precise interplay between these factors and their combined contribution to overall capacity remains not fully understood. We explore this relationship in the context of sparse Mixture-of-Expert models (MoEs), which allow scaling the number of parameters without proportionally increasing the FLOPs per example. We investigate how varying the sparsity level, i.e., the ratio of non-active to total parameters, affects model performance in terms of both pretraining and downstream performance. We find that under different constraints (e.g. parameter size and total training compute), there is an optimal level of sparsity that improves both training efficiency and model performance. These results provide a better understanding of the impact of sparsity in scaling laws for MoEs and complement existing works in this area, offering insights for designing more efficient architectures.
@@ 1 Introduction
Our results, inline with findings from previous relevant studies (ludziejewski2024scaling; he2024mixture) on scaling laws for MoEs, show increasing sparsity level leads to better performance and efficiency during pretraining. Considering the various methods to increase compute per example during inference adaptively conditioned on task or example complexity, we conclude that approaches like Mixture-of-Experts (MoEs), which reduce the unit compute cost (i.e., FLOPs per token) by increasing the sparsity level, hold significant promise given their potential to enhance efficiency in both pretraining and inference.
@@
Figure 1: IsoFLOP surface over observed pretraining loss $\mathbf{L}$, model size (in terms of total $\mathbf{N}$ and active parameters $\mathbf{N_{a}}$), and sparsity $\mathbf{S}$. We fit a polynomial function mapping $\mathbf{N}$ (or $\mathbf{N_{a}}$), $\mathbf{S}$, and their interaction to $\mathbf{L}$, using empirical data. For both fits the MSE loss for predicting loss on a held out set is $0.0001$. These results indicate that for a fixed compute budget, increasing model sparsity leads to a reduction in pretraining loss. When considering optimal model size, we observe opposite trends for total parameters ($\mathbf{N}$) (Figure a) versus active parameters ($\mathbf{N_{a}}$) (Figure b). (See Figure 8 in Section D.1 for results with different total compute budgets $C$.)
@@ 2 The Interplay between Model Parameters and Sparsity in MoEs
As the first step, considering a fixed training compute budget $C$, we fit a 3D surface, referred to as the IsoFLOP surface, in Figure 1a, using a polynomial function, following approach II of hoffmann2022training. Compared to hoffmann2022training we include the sparsity variable and fit a single 3d IsoFLOP surface across all data points, rather than fitting separate 2d IsoFLOP curves for fixed sparsity levels or model sizes. We conducted a grid search to determine the optimal polynomial degree for $N$, $S$, and the interaction term $N\times S$, finding that a degree of $(2,2,2)$ resulted in the lowest cross-validation error. Both $N$ and $S$ are in log space (see Appendix B for more details). Figure 1b illustrates the goodness of fit, demonstrating a strong correlation and low predictive error.
@@ 2.2 Optimal Sparsity Level for Fixed Model Size
Figure 3: Effect of compute budget on model size, number of active parameters and loss with sparsity. Over all budgets considered, we observe that (a) the optimal model size $N$ increases with sparsity, (b) the optimal number of active parameters $N_{a}$ decreases with sparsity, and (c) the loss $L$ decreases with sparsity.
@@ 3 Impact of Training Compute Budget on the Interaction between Model Parameters and Sparsity
Does increasing compute budget impact the interaction between the parameters and compute per example in MoEs and how they contribute to model’s capacity? In other words, does the recipe for optimally increasing model capacity, i.e., optimal sparsity level for MoEs change as we scale up the total training compute?
@@
Figure 4: Effect of training budget $C$ and total parameters $N$ on MoE sparsity. This figure shows how the optimal MoE sparsity $S^{*}$ changes with respect to the total number of parameters $N$ and the training budget $C$. The $x$-axis represents the total parameters $N$ on a logarithmic scale, and the $y$-axis shows the optimal MoE sparsity $S^{*}$.
@@
In Figures 3a and 3b, , we see a consistent trend of increasing $N$ and decreasing $N_{a}$ for compute optimal models as sparsity level increases across all training compute budgets. Moreover, as can be seen in Figure 4, when model size in terms of total number of parameters is fixed, optimal sparsity level increases with training compute budget as well as model size as discussed in Section 2.2.
@@ 4 Effect of MoE Sparsity on Downstream Task Performance
We use downstream tasks from the evaluation suite in llm-foundry [^3] Github repository: https://github.com/mosaicml/llm-foundry for benchmarking our pretrained models. The downstream task are devided into four pre-defined categories namely: language understanding, world knowledge, reading comprehension, and symbolic reasoning to help us systematically test whether the downstream vs upstream performance trend remains the same or is different as we vary sparsity values.
@@
We observe from  Figure 5a (language understanding),  Figure 5c (commonsense reasoning), and  Figure 5d (world knowledge) that there is a strong correlation between upstream (pretraining) loss and downstream performance (error) across all these tasks. For these tasks, downstream performance is predictable based on upstream performance, regardless of the sparsity level. This indicates that, in the context of these tasks, the optimal sparsity level follows the same trend as the optimal sparsity observed during pretraining. However, Figure 5b (reading comprehension) shows an example of a task where models with higher sparsity transfer more poorly compared to denser models. This decrease in the transfer performance of sparser models on these tasks may be due to the lower inference-time compute in sparser models compared to their denser counterparts for a similar pretraining loss. Further analysis is needed to verify this intuition.
If fewer FLOPs per example are the reason behind the worse transfer performance in sparser models, this effect might diminish with a larger total training compute budget, as the optimal active number of parameters increases. Moreover, one can use approaches like chain-of-thought reasoning to independently increase FLOPs per example during inference time. In Appendix E, we explore whether increasing inference-time compute via Chain-of-Thought (CoT) prompting can improve the performance of MoEs on tasks that require more reasoning. Our experiments indicate that MoEs benefit more from this increased compute compared to dense models with a similar number of active parameters. This suggests that dynamic compute allocation during inference may be crucial for MoEs to perform well on complex reasoning tasks.
@@ 5 Incorporating Sparsity into Scaling Laws
For dense models with a fixed total training FLOPs, $C$, the parameters $N$ and $D$ are interrelated through the equation for estimating FLOPs per example, given as $C=6ND$ for transformers. However, in MoEs (Mixture of Experts models), this relationship involves the active number of parameters $N_{a}$ rather than the total parameter count $N$. Thus, $D$ and $N_{a}$ define the total training FLOPs rather than $D$ and $N$. Given the analysis conducted in §2, we know that if the total number of parameters $N$ is fixed, the optimal sparsity level, i.e., active number of parameters would depend on $N$. Motivated by this observation, we suggest the following parametric form that includes a a multiplicative interaction between $N$ and $S$ or $N_{a}$ to predict the loss:
@@
By incorporating sparsity into the scaling law equation, we can eliminate the need for parameters specific to mixture-of-experts (MoE) models, such as the total and active number of experts. Indeed, as frantar2024scaling demonstrates, this same equation continues to hold under a different mechanism for introducing sparsity—weight sparsity—where individual neural network connections are pruned.
We use the recipe described by hoffmann2022training and use the L-BFGS algorithm to fit the coefficients in Equation 6 using a Huber loss with $\delta=10^{-3}$. The optimal coefficient values are found via a grid search with values described in Table 2. The results of data fitting and validation are shown in Figure 6. The estimated values are shown in Table 3 in Appendix F.
@@
Figure 6: Scaling law fit on data obtained from training compute-optimal models. Figure (a) shows fit on data used to estimate coefficients for Equation 6 while (b) validates the estimated coefficients on a held-out dataset. All $S=0.98$ data points were excluded from the fitting process to provide an out-of-sample validation.
@@ 6 Discussion
  • Effect of Sparsity on Downstream Performance: For most downstream tasks, models with similar pretraining perplexity have similar downstream task performance regardless of sparsity. For reading comprehension tasks (e.g., CoQA (reddy-etal-2019-coqa), SQuAD (rajpurkar-etal-2018-know)), denser models perform better, potentially due to their higher inference-time compute than a perplexity-matched sparse model. Strategies to increase inference time compute dynamically (wei2022chain_cot; goyal2024think_pause) may address this gap. We use a “length-controlled” variant of few-shot Chain-of-Thought (CoT) prompting to study whether this strategy may help close the gap between MoEs and dense models on reasoning tasks in Appendix E. Empirical evidence suggests that MoEs may benefit more from dynamically increased inference-time compute than dense models under a fixed generated tokens budget.
  • Parametric Scaling Law: We propose a parametric form for scaling laws that accounts for sparsity. The model coefficients are estimated using the empirical data obtained by training compute-optimal models. An interesting observation from  Appendix F is that the exponent for sparsity term $\lambda$ is negative which is consistent with our intuition that sparser models lead to a lower perplexity.
@@ 8 Conclusion
In this paper, we try to understand the optimal trade-off between parameters and compute per example to optimally increase the model capacity. Our experiments and analysis, underscores the role of sparsity, as a factor that controls the balance between total number of parameters and FLOPs per example, in the scaling laws for Mixture-of-Expert Transformers (MoEs), showing that the most efficient model configuration depends on balancing model size, training compute, and sparsity level. The optimal recipe for balancing FLOPs per example and parameter count in MoEs depends on the objective as well other resource constraints. Our findings indicate that sparsity, as a knob that controls FLOPs per example in MoEs, is a powerful mechanism for optimizing model performance under constrained training compute budgets. By balancing the total number of parameters, compute, and sparsity, MoEs can be scaled more effectively. These insights provide valuable guidance for scaling language models, especially for MoEs, where the trade-offs between parameters and FLOPs must be carefully managed.
MoEs were originally introduced to allow increasing model capacity without a significant increase in inference cost. Our experiments show that under fixed total training compute budget increasing sparsity in MoEs leads to smaller FLOPs per example, higher number of parameters, and lower pretraining loss simultaneously. In other words, in the context of MoEs, if there are no constraints on the total number of parameters, increasing the capacity of the model through parameter count seem to be the optimal strategy if lower pretraining loss is the main goal. On the other hand, when comparing how well the pretraining performance transfers to various downstream tasks, denser models seem to be better on certain types of task that potentially rely on deeper processing of the input vs the knowledge stored in the parameters of the model. This potentially signals the importance of the role of FLOPs per example in increasing the capacity of the model during inference. Our experiments demonstrate that Mixture-of-Experts (MoE) models use Chain-of-Thought (CoT) prompting more effectively than dense models, achieving better performance when allocated additional computational resources during inference. This observation reveals an interesting direction to improve the performance efficiency of MoEs at inference time.
@@
Another important avenue will be to examine how the findings on the role of sparsity in Mixture-of-Experts (MoEs) generalize to architectures or approaches that employ different mechanisms for independently adjusting FLOPs per example and the number of trainable parameters. Additionally, an intriguing direction for future exploration is the study of scaling behaviors in models that enable negative sparsity values through parameter sharing.
@@ Appendix A Preliminaries
In this section, we provide a brief overview of Mixture-of-Expert (MoE) Transformers.
A.1 Mixture-of-Expert (MoE) Transformers
@@ D.4 Comparing IsoFLOP Surface Analysis with Independent 2d IsoFLOPs
Appendix E Does Chain-of-Thought prompting benefit sparse MoEs more than dense models?
In Section 4, we observed that dense models fare marginally better than sparse MoEs on reading comprehension tasks, potentially due to the higher inference-time compute of a dense model than a perplexity-matched sparse MoE. Then, in Section 6, we hypothesized that alternative strategies to increase inference-time compute may reduce the gap between sparse MoEs and dense models on such tasks. In this section, we test this hypothesis by leveraging a “length-controlled” variant of few-shot Chain-of-Thought (CoT) prompting to indirectly control inference-time compute. We then use this to study the effect of inference-time compute on downstream task performance of dense and MoE models.
Experiment setup.
We evaluate Qwen1.5 models (qwen) on the GSM8k dataset (cobbe2021gsm8k) to study the effect of few-shot CoT prompting (wei2022chain_cot) on downstream task performance. We look at the effect of increasing inference-time compute (via CoT prompting) on GSM8k performance of dense models with sizes ranging from 0.5B to 14B and a 5x2.7B sparse MoE. We also use $10$ fixed examples from the GSM8k train split as few-shot examples for all runs.
Length-controlled CoT prompting enables control over inference-time compute.
Given an instruction-tuned model and a problem from the GSM8k dataset, we control the inference-time compute of the model by controlling the number of tokens generated to output the final answer to the given problem. To do so, we observe that providing instructions via system prompts (with few-shot CoT prompting) does not effectively control the number of generated tokens and, as a result, inference-time compute. We also observe that the average number of tokens in the few-shot GSM8k answers (provided in-context) strongly influences the number of tokens generated by the model to solve the given GSM8k question. Therefore, similar to work on designing GSM8k variants to analyze language modeling phenomena (mirzadeh2024gsm; li2024gsmplus; zhang2024careful), we prompt an instruction-tuned model—Llama-3.1-70b (dubey2024llama) in our experiment—to rewrite GSM8k answers in approximately $k\in\{5,\dots,100\}$ words. Then, we use the paraphrased GSM8k examples as few-shot examples to indirectly control the number of generated tokens. As shown in Figure 13(a), this approach enables systematic control over the length of paraphrased answers. The strong correlation ($\rho=0.88$) between few-shot and generated answer lengths in Figure 13(b) validates that our approach effectively modulates inference-time compute.
Figure 13: Varying inference-time compute via length-controlled Chain-of-Thought prompting. We can control inference-time compute (number of generated tokens) via Chain-of-Thought prompting in two steps: (a) generating paraphrased GSM8k answers of varying lengths (5-100 words) and (b) using paraphrased answers as few-shot examples in CoT prompts to influence the length of generated answers determine the model’s output length ($\rho=0.88$ correlation). The systematic shift in answer length distributions in subplot (a) and the linear relationship between few-shot answer length and generated answer lenth ($\rho=0.88$ correlation) in (b) validate that our prompting approach effectively modulates inference-time compute, enabling controlled studies of its impact on model performance.
Effect of length-controlled CoT on downstream task performance.
We investigate how inference-time compute affects GSM8k performance across Qwen1.5 models using $10$-shot length-controlled CoT prompting, varying target answer lengths $k$ from $1$ to roughly $70$ words on average. As shown in Figure 14, when we increase inference-time compute indirectly through longer generated answers ($x$-axis), we observe a roughly linear improvement in GSM8k performance ($y$-axis) across all model sizes. The linear fits for each model also indicate a pattern through their slopes $m$: larger dense models benefit more from additional inference-time compute. This effect is particularly striking when comparing models at the extremes (i.e., Qwen1.5-0.5B versus Qwen1.5-14B).
To analyze whether inference-time compute affects dense and sparse MoE models differently, we examine how the “performance” slope $m$ (i.e., accuracy gain per generated word) varies with model size. While the Qwen1.5-5x2.7B MoE cannot be directly compared to dense models due to different active parameter counts, we can account for this by plotting $m$ against the number of active parameters for both architectures. As shown in Figure 15, when controlling for active parameter count, the MoE model exhibits a higher slope than would be expected from interpolating between dense models. This suggests that MoE models benefit more from dynamically increased inference-time compute compared to dense models with equivalent active parameters.
Figure 14: Effect of length-controlled CoT prompting on GSM8k performance across model scales. We evaluate the relationship between inference-time compute (controlled via answer length) and GSM8k accuracy for dense Qwen1.5 models (0.5B-14B parameters) and a 5x2.7B sparse MoE. For all models, increased inference-time compute improves accuracy roughly linearly, with slopes $m$ indicating the marginal effect. Larger dense models show steeper slopes, demonstrating they benefit more from additional inference-time compute—for example, the 14B model’s performance improves from 20% to 70% as answer length increases, while the 0.5B model improves only from 1% to 5%.
Figure 15: Sparse MoEs benefit more from increased inference-time compute than dense models. We plot the marginal effect of inference-time compute (accuracy gain per generated word) against model size for dense Qwen1.5 models and a 5x2.7B sparse Qwen1.5 MoE. When controlling for the number of active parameters, the MoE model (orange) shows a higher performance slope than would be expected from interpolating between dense models (blue), suggesting that sparse MoEs benefit more from dynamically increased inference-time compute via strategies like CoT prompting.
Appendix F Incorporating Sparsity into Scaling Laws

### v2

@@ Abstract
Scaling the capacity of language models has consistently proven to be a reliable approach for improving performance and unlocking new capabilities. Capacity can be primarily defined by two dimensions: the number of model parameters and the compute per example. While scaling typically involves increasing both, the precise interplay between these factors and their combined contribution to overall capacity remains not fully understood. We explore this relationship in the context of sparse Mixture-of-Experts (MoEs) , which allow scaling the number of parameters without proportionally increasing the FLOPs per example. We investigate how varying the sparsity level, i.e., the fraction of inactive parameters, impacts model’s performance during pretraining and downstream few-shot evaluation. We find that under different constraints (e.g., parameter size and total training compute), there is an optimal level of sparsity that improves both training efficiency and model performance. These results provide a better understanding of the impact of sparsity in scaling laws for MoEs and complement existing works in this area, offering insights for designing more efficient architectures.
@@ 1 Introduction
Our results, in line with findings from previous relevant studies (ludziejewski2024scaling; he2024mixture) on scaling laws for MoEs, show increasing sparsity level leads to better performance and efficiency during pretraining. Considering the various methods to increase compute per example during inference adaptively conditioned on task or example complexity, we conclude that MoEs approaches, which reduce the unit compute cost (i.e., FLOPs per token) by increasing the sparsity level, hold significant promise given their potential to enhance efficiency in both pretraining and inference.
@@
Figure 1: IsoFLOP surface over observed pretraining loss $\mathbf{L}$, model size (in terms of total $\mathbf{N}$ and active parameters $\mathbf{N_{a}}$), and sparsity $\mathbf{S}$. We fit a polynomial function mapping $\mathbf{N}$ (or $\mathbf{N_{a}}$), $\mathbf{S}$, and their interaction to $\mathbf{L}$, using empirical data. Both fits achieve an MSE loss of 0.0001 on a held-out set. These results indicate that for a fixed compute budget, increasing model sparsity leads to a reduction in pretraining loss. When considering optimal model size, we observe opposite trends for total parameters ($\mathbf{N}$) (Figure a) versus active parameters ($\mathbf{N_{a}}$) (Figure b). (Figure 8 includes results Section D.1 for different total compute budgets $C$.)
@@ 2 The Interplay between Model Parameters and Sparsity in MoEs
As the first step, considering a fixed training compute budget $C$, we fit a 3D surface, referred to as the IsoFLOP surface, in Figure 1a, using a polynomial function, following approach II of hoffmann2022training. Compared to hoffmann2022training we include the sparsity variable and fit a single 3d IsoFLOP surface across all data points, rather than fitting separate 2d IsoFLOP curves for fixed sparsity levels or model sizes. We conducted a grid search to determine the optimal polynomial degree for $N$, $S$, and the interaction term $N\times S$, finding that a degree of $(2,2,2)$ resulted in the lowest cross-validation error. Both $N$ and $S$ are in log space (see Appendix B for more details).
@@ 2.2 Optimal Sparsity Level for Fixed Model Size
Figure 3: Effect of compute budget on model size, number of active parameters and loss with sparsity. Across all compute budgets, we observe that (a) the optimal model size $N$ increases with sparsity, (b) the optimal number of active parameters $N_{a}$ decreases with sparsity, and (c) the loss $L$ decreases with sparsity.
@@ 3 Impact of Training Compute Budget on the Interaction between Model Parameters and Sparsity
Does increasing compute budget impact the interaction between the parameters and FLOPs per example in MoEs and how they contribute to model’s capacity? In other words, does the recipe for optimally increasing model capacity, i.e., optimal sparsity level for MoEs change as we scale up the total training compute?
@@
Figure 4: Effect of training budget $C$ and total parameters $N$ on MoE sparsity. Optimal MoE sparsity $S^{*}$ changes with respect to the total number of parameters $N$ and the training budget $C$. The $x$-axis represents the total parameters $N$ on a logarithmic scale, and the $y$-axis shows the optimal MoE sparsity $S^{*}$.
@@
In Figure 3a and Figure 3b, , we see a consistent trend of increasing $N$ and decreasing $N_{a}$ for compute optimal models as sparsity level increases across all training compute budgets. Moreover, as can be seen in Figure 4, when model size in terms of total number of parameters is fixed, optimal sparsity level increases with training compute budget as well as model size as discussed in Section 2.2.
@@ 4 Effect of MoE Sparsity on Downstream Task Performance
We use downstream tasks from the evaluation suite in llm-foundry [^3] Github repository: https://github.com/mosaicml/llm-foundry for benchmarking our pretrained models, specifically in an in-context few-shot learning setup. This setup focuses on evaluating a model’s ability to learn and adapt to new tasks with limited examples. The downstream task are devided into four pre-defined categories namely: language understanding, world knowledge, reading comprehension, and symbolic reasoning to help us systematically test whether the downstream vs upstream performance trend remains the same or is different as we vary sparsity values.
@@
We observe from  Figure 5a (language understanding),  Figure 5c (commonsense reasoning), and  Figure 5d (world knowledge) that, in an in-context few-shot learning setting, there is a strong correlation between upstream (pretraining) loss and downstream performance (error) across all these tasks. For these tasks, downstream performance in the few-shot setting is predictable based on upstream performance, regardless of the sparsity level. This indicates that, in the context of these tasks, the optimal sparsity level follows the same trend as the optimal sparsity observed during pretraining. However, Figure 5b (reading comprehension) shows an example of a task where models with higher sparsity transfer more poorly compared to denser models. This decrease in the transfer performance of sparser models on these tasks may be due to the lower inference-time compute in sparser models compared to their denser counterparts for a similar pretraining loss. Further analysis is needed to verify this intuition.
If fewer FLOPs per example are the reason behind the worse transfer performance in sparser models, this effect might diminish with a larger total training compute budget, as the optimal active number of parameters increases. Moreover, one can use approaches like chain-of-thought reasoning (wei2022chain_cot) to independently increase FLOPs per example during inference time.
@@ 5 Incorporating Sparsity into Scaling Laws
For dense models with a fixed total training FLOPs, $C$, the parameters $N$ and $D$ are interrelated through the equation for estimating FLOPs per example, given as $C=6ND$ for transformers. However, in MoEs (Mixture of Experts models), this relationship involves the active number of parameters $N_{a}$ rather than the total parameter count $N$. Thus, $D$ and $N_{a}$ define the total training FLOPs rather than $D$ and $N$. Given the analysis conducted in §2, we know that if the total number of parameters $N$ is fixed, the optimal sparsity level, i.e., active number of parameters would depend on $N$. Motivated by this observation, we suggest the following parametric form that includes a multiplicative interaction between $N$ and $S$ or $N_{a}$ to predict the loss:
@@
By incorporating sparsity into the scaling law equation, we can eliminate the need for parameters specific to MoEs, such as the total and active number of experts. As demonstrated by frantar2024scaling, this formulation also holds for other sparsity mechanisms, such as weight sparsity, where individual neural network connections are pruned.
We use the recipe described by hoffmann2022training and use the L-BFGS algorithm to fit the coefficients in equation 6 using a Huber loss with $\delta=10^{-3}$. Optimal coefficient values were determined through a grid search (see Table 2 for search values). The results of data fitting and validation are shown in Figure 6. The estimated values are shown in Table 3 in Appendix E.
@@
Figure 6: Scaling law fit on data obtained from training compute-optimal models. Figure 6(a) shows the fit on the data used to estimate the coefficients for equation 6, while Figure 6(b) validates these coefficients on a held-out dataset. All data points with $S=0.98$ were excluded from the fitting process for out-of-sample validation.
@@ 6 Discussion
  • Effect of Sparsity on Downstream Performance: For most downstream tasks, models with similar pretraining perplexity have similar downstream task performance regardless of sparsity. For reading comprehension tasks (e.g., CoQA (reddy-etal-2019-coqa), SQuAD (rajpurkar-etal-2018-know)), denser models perform better, potentially due to their higher inference-time compute than a perplexity-matched sparse model. Strategies to increase inference time compute dynamically (wei2022chain_cot; goyal2024think_pause) may address this gap.
  • Parametric Scaling Law: We propose a parametric form for scaling laws that accounts for sparsity. The model coefficients are estimated using the empirical data obtained by training compute-optimal models. An interesting observation from  Appendix E is that the exponent for sparsity term $\lambda$ is negative which is consistent with our intuition that sparser models lead to a lower perplexity.
@@ 8 Conclusion
In this paper, we investigated the optimal trade-off between parameters and compute per example for maximizing model capacity. Our findings indicate that sparsity, as a knob that controls FLOPs per example in MoEs, is a powerful mechanism for optimizing model performance under constrained training compute budgets. By balancing the total number of parameters, compute, and sparsity, MoEs can be scaled more effectively. These insights provide valuable guidance for scaling language models, especially for MoEs, where the trade-offs between parameters and FLOPs must be carefully managed.
MoEs were originally introduced to allow increasing model capacity without a significant increase in inference cost. Our experiments show that under fixed total training compute budget increasing sparsity in MoEs leads to smaller FLOPs per example, higher number of parameters, and lower pretraining loss simultaneously. In other words, in the context of MoEs, if there are no constraints on the total number of parameters, increasing the capacity of the model through parameter count seem to be the optimal strategy if lower pretraining loss is the main goal. On the other hand, when comparing how well the pretraining performance transfers to various downstream tasks, denser models exhibit better transfer performance on certain types of task that potentially rely on deeper processing of the input vs the knowledge stored in the parameters of the model. This potentially signals the importance of the role of FLOPs per example in increasing the capacity of the model during inference. This observation reveals an interesting direction to improve the performance efficiency of MoEs at inference time.
@@
Another important avenue will be to examine how the findings on the role of sparsity in MoEs generalize to architectures or approaches that employ different mechanisms for independently adjusting FLOPs per example and the number of trainable parameters. Additionally, an intriguing direction for future exploration is the study of scaling behaviors in models that enable negative sparsity values through parameter sharing.
@@ Appendix A Preliminaries
A.1 Notation and Terminology
To aid readability, we provide a list of key symbols used throughout this paper.
Symbol Description
$N$ Total number of model parameters
$N_{a}$ Active number of model parameters
$S$ Sparsity level (ratio of non-active to total experts)
$S^{*}$ Optimal sparsity level
$L$ Pretraining Loss (Categorical Cross-Entropy)
$L^{*}$ Optimal pretraining loss
$C$ Total training compute budget (in FLOPs)
$N^{*}$ Optimal total number of parameters
$N^{*}_{a}$ Optimal active number of parameters
$E$ Expansion factor (number of experts per MoE layer)
$K$ Number of selected experts per token
$G$ Granularity of experts (size relative to base MLP)
$D$ Dataset size (number of training tokens)
$\alpha,\beta,\gamma,\lambda,\delta,a,b,c,d,e$ Coefficients in the parametric scaling law equation
In this paper, we use the term "compute" in a general sense to refer to computational cost. Unless otherwise specified, "compute" and "FLOPs" (Floating Point Operations) are used interchangeably to quantify this cost.
A.2 Mixture-of-Expert (MoE) Transformers
@@ D.4 Comparing IsoFLOP Surface Analysis with Independent 2d IsoFLOPs
Appendix E Incorporating Sparsity into Scaling Laws